# DeRO: Dead Reckoning Based on Radar Odometry With Accelerometers Aided for Robot Localization

### v2

@@ III-B System Overview
The overview of the proposed DeRO is depicted in Fig. 2. In this study, we apply a stochastic cloning based error-state (indirect) EKF (SC-IEKF) to fuse data from an IMU and a 4D FMCW radar. The raw point clouds obtained from radar are first fed into an outlier rejection algorithm. Here, we employ a simple RANSAC-based least square (RANSAC/LSQ) method introduced in [11]. The outputs of this algorithm include the estimated ego velocity $\hat{v}^{r}$ along with its corresponding error covariance $\hat{Q}$, as well as a set of inlier point clouds. These velocity estimates and covariance matrices are used for radar mechanization in conjunction with gyroscope data to compute poses and the filter’s time update. Subsequently, the inlier point cloud set is merged with the inlier set from the one at $M$ preceding steps to perform scan matching to calculate the relative distance estimate $\hat{p}^{r_{k}}_{r_{k-M}}$ with error covariance $\hat{R}$ between these two radar scans. This information will update the estimation.
@@
III-C Radar mechanization
@@ IV-B System Model
where $T$ represents the radar sampling period, and the dynamic matrix $F\in\mathbb{R}^{12\times 12}$ is defined (for brevity, only non-zero elements of $F$ are provided2)
@@ IV-C1 Radar range measurements
Given the relative distance $\bar{p}_{r_{k}}^{r_{k-M}}$ between the step $k$ and $k-M$ of the radar frame obtained from a scan matching algorithm. By solving a straightforward vector geometry problem, one could establish the following relationship
$\hat{p}_{r_{k}}^{r_{k-M}}$ $=C_{b}^{r}\hat{C}_{n}^{b_{k-M}}\big{(}\hat{p}_{b_{k}}^{n}-\hat{p}_{b_{k-M}}^{n}\big{)}.$ (21)
@@
$\breve{z}_{r,k}$ $=\hat{p}_{r_{k}}^{r_{k-M}}-\bar{p}_{r_{k}}^{r_{k-M}}=\breve{H}_{r,k}\delta x_{k}+n_{p_{r}},$ (22)
@@
$\breve{H}_{r,k}$ $=C_{b}^{r}\hat{C}_{n}^{b_{k-M}}\begin{bmatrix}I_{3}&\left[\hat{C}_{b_{k}}^{n}p_{r}^{b}\right]_{\!\times}&0_{3\times 6}&-I_{3}&\Lambda\end{bmatrix},$ (23)
$\Lambda=\left[\hat{p}_{b_{k-M}}^{n}-\hat{p}_{b_{k}}^{n}-\hat{C}_{b_{k}}^{n}p_{r}^{b}\right]_{\!\times}$, and $n_{p_{r}}\thicksim{\mathcal{N}}\big{(}0_{3\times 1},\sigma^{2}_{p_{r}}I_{3}\big{)}$.
@@ IV-C2 Tilt angles
$z_{a,k}$ $=\begin{bmatrix}\hat{\phi}\\ \hat{\theta}\end{bmatrix}-\begin{bmatrix}\bar{\phi}_{a}\\ \bar{\theta}_{a}\end{bmatrix}=H_{a,k}\delta\Psi_{b,k}^{n}+n_{a}$
$=\begin{bmatrix}-\dfrac{\cos(\hat{\psi})}{\cos(\hat{\theta})}&-\dfrac{\cos(\hat{\psi})}{\cos(\hat{\theta})}&0\\[8.00003pt] \sin(\hat{\psi})&-\cos(\hat{\psi})&0\end{bmatrix}\delta\Psi_{b,k}^{n}+n_{a}.$ (25)
Here, the estimate Euler angle $\hat{\Psi}_{b}^{n}=\begin{bmatrix}\hat{\phi}&\hat{\theta}&\hat{\psi}\end{bmatrix}^{\top}$ can be found by transforming the estimate quaternion $\hat{q}_{b}^{n}$ and $n_{a}\thicksim{\mathcal{N}}\big{(}0_{2\times 1},\sigma^{2}_{a}I_{2}\big{)}$. Furthermore, given that velocity information is available from radar measurements, we compensate for the linear acceleration and biases derived from accelerometer output. This compensation enhances the quality of tilt angle information. Particularly,
$\hat{f}^{b}=\bar{f}^{b}-\hat{b}_{a}-\hat{C}_{n}^{b}\dfrac{\big{(}\hat{v}_{b_{k}}^{n}-\hat{v}_{b_{k-1}}^{n}\big{)}}{T}.$ (26)
@@ IV-D Implementation
12:         Radar mechanization.
@@
15:              Registration $\rightarrow\hat{p}_{r_{k}}^{r_{k-M}},\hat{\sigma}_{p_{r}}^{2}$.

### v3

@@ III-B System Overview
The overview of the proposed DeRO is depicted in Fig. 2. In this study, we apply a stochastic cloning based error-state (indirect) EKF (SC-IEKF) to fuse data from an IMU and a 4D FMCW radar. The raw point clouds obtained from radar are first fed into an outlier rejection algorithm. Here, we employ a simple RANSAC-based least square (RANSAC/LSQ) method introduced in [11]. The outputs of this algorithm include the estimated ego velocity $\hat{v}^{r}$ along with its corresponding error covariance $\hat{Q}$, as well as a set of inlier point clouds. These velocity estimates and covariance matrices are used for radar-based DR mechanization in conjunction with gyroscope data to compute poses and the filter’s time update. Subsequently, the inlier point cloud set is merged with the inlier set from the one at $M$ preceding steps to perform scan matching to calculate the relative distance estimate $\hat{p}^{r,k}_{r,{k-M}}$ with error covariance $\hat{R}$ between these two radar scans. This information will update the estimation.
@@
III-C Radar-based Dead Reckoning Mechanization
@@ IV-B System Model
where $T$ denotes the radar sampling period, and the continuous-time error dynamics matrix $F\in\mathbb{R}^{12\times 12}$ is given (for brevity, only non-zero elements of $F$ are provided2)
@@ IV-C1 Radar range measurements
Given the relative distance $\bar{p}_{r,k}^{r,{k-M}}$ between the step $k$ and $k-M$ of the radar frame obtained from a scan matching algorithm. By solving a straightforward vector geometry problem, one could establish the following relationship
$\hat{p}_{r,k}^{r,{k-M}}$ $=C_{b}^{r}\hat{C}_{n}^{b,{k-M}}\big{(}\hat{p}_{b,{k}}^{n}-\hat{p}_{b,{k-M}}^{n}\big{)}.$ (21)
@@
$\breve{z}_{r,k}$ $=\hat{p}_{r,k}^{r,{k-M}}-\bar{p}_{r,k}^{r,{k-M}}=\breve{H}_{r,k}\delta x_{k}+n_{p_{r}},$ (22)
@@
$\breve{H}_{r,k}$ $=C_{b}^{r}\hat{C}_{n}^{b,{k-M}}\begin{bmatrix}I_{3}&\left[\hat{C}_{b,k}^{n}p_{r}^{b}\right]_{\!\times}&0_{3\times 6}&-I_{3}&\Lambda\end{bmatrix},$ (23)
$\Lambda\!=\!\left[\hat{p}_{b,{k-M}}^{n}-\hat{p}_{b,{k}}^{n}-\hat{C}_{b,k}^{n}p_{r}^{b}\right]_{\!\times}$, and $n_{p_{r}}\!\!\thicksim{\mathcal{N}}\big{(}0_{3\times 1},\sigma^{2}_{p_{r}}I_{3}\big{)}$.
@@ IV-C2 Tilt angles
$z_{a,k}$ $=\begin{bmatrix}\hat{\phi}_{k}\\ \hat{\theta}_{k}\end{bmatrix}-\begin{bmatrix}\bar{\phi}_{a,k}\\ \bar{\theta}_{a,k}\end{bmatrix}=H_{a,k}\delta\Psi_{b,k}^{n}+n_{a}$
$=\begin{bmatrix}-\dfrac{\cos(\hat{\psi}_{k})}{\cos(\hat{\theta}_{k})}&-\dfrac{\cos(\hat{\psi}_{k})}{\cos(\hat{\theta}_{k})}&0\\[8.00003pt] \sin(\hat{\psi}_{k})&-\cos(\hat{\psi}_{k})&0\end{bmatrix}\delta\Psi_{b,k}^{n}+n_{a}.$ (25)
Here, the estimate Euler angle $\hat{\Psi}_{b,k}^{n}=\begin{bmatrix}\hat{\phi}_{k}&\hat{\theta}_{k}&\hat{\psi}_{k}\end{bmatrix}^{\top}$ can be found by transforming the estimate quaternion $\hat{q}_{b,k}^{n}$ and $n_{a}\thicksim{\mathcal{N}}\big{(}0_{2\times 1},\sigma^{2}_{a}I_{2}\big{)}$. Furthermore, given that velocity information is available from radar measurements, we compensate for the linear acceleration and biases derived from accelerometer output. This compensation enhances the quality of tilt angle information. Particularly,
$\hat{f}^{b}_{k}=\bar{f}^{b}_{k}-\hat{b}_{a,k}-\hat{C}_{n}^{b,k}\dfrac{\big{(}\hat{v}_{b,k}^{n}-\hat{v}_{b,{k-1}}^{n}\big{)}}{T}.$ (26)
@@ IV-D Implementation
12:         Radar-based dead reckoning mechanization.
@@
15:              Registration $\rightarrow\hat{p}_{r,k}^{r,{k-M}},\hat{\sigma}_{p_{r}}^{2}$.